please answer these2. Because many passengers who make reservations do not show up, airlines often overbook flights (sell more tickets than ther questions
Answer #1

a.

Number of tickets sold, n = 255

Probability that a passenger show up = 1 - 0.05 = 0.95

Using Normal approximation, the number of passengers showing up will follow Normal distribution with mean = np = 255 * 0.95 = 242.25 and standard deviation = Vnp(1 - p) = 255 * 0.95 * (1 -0.95) = 3.48

Probability that at least 246 showing up = P(X \ge 246) = P(X > 245.5) (Using Continuity correction)

= P[Z > (245.5 - 242.25) / 3.48]

= P[Z > 0.9339]

= 0.1752

b.

Since the probability that at least 246 showing up is not very low, (not less than 0.05), the airline should change the number of tickets it sells.

3.

Sample proportion, \hat{p} = 66/200 = 0.33

Hypothesized proportion, p = 0.30

If president's skepticism is justified, the standard error of sample proportion is,

SE = p(1 - p)/n= 0.3* (1 -0.3)/200 = 0.0324

Probability of finding a sample proportion equal to or greater than that found if the president's skepticism is justified

= P(p \ge 0.33 | p = 0.30)

= P[Z \ge (0.33 - 0.30) / 0.0324]

= P[Z \ge 0.9259]

= 0.1772